# Multiple quantum exceptional, diabolical, and hybrid points in multimode bosonic systems: I. Inherited and genuine singularities

### v1

@@ V.2 Pyramid configuration
$\gamma_{+}$ $-\zeta\mp\beta_{2}$ $\langle b_{4}\rangle$, $\langle b_{5}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{1}}\rangle$ 1 1x2 1x2 1x2 1x2
$\zeta\pm\beta_{2}$ $\langle b_{4}^{\dagger}\rangle$, $\langle b_{5}\rangle$ $\langle\bm{\hat{B}_{2}}\rangle$ 1 1x2 1x2 1x2 1x2
$\zeta-\beta_{1}$ $\langle b_{2}\rangle$ $\langle\bm{\hat{B}_{3}}\rangle$ 1 1x1 1x1 1x1 1x1
$-\zeta+\beta_{1}$ $\langle b_{2}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{4}}\rangle$ 1 1x1 1x1 1x1 1x1
$-\zeta-\beta_{1}$ $\langle b_{3}\rangle$ $\langle\bm{\hat{B}_{5}}\rangle$ 1 1x1 1x1 1x1 1x1
$\zeta+\beta_{1}$ $\langle b_{3}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{6}}\rangle$ 1 1x1 1x1 1x1 1x1
0 $\langle b_{1}\rangle$ $\langle\bm{\hat{B}_{7}}\rangle$ 1 2x1 2x1 2x1 2x1
$\langle b_{1}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{8}}\rangle$ 1
$2\gamma_{+}$ $-2\zeta\mp 2\beta_{2}$ $\langle b_{4}^{2}\rangle$, $\langle b_{5}^{\dagger 2}\rangle$ $\langle\bm{\hat{B}_{1}^{2}}\rangle$ 1 1x4 1x4 1x3 1x3
$-2\zeta$ $\langle b_{5}^{\dagger}b_{4}\rangle$ 2
$2\zeta\pm 2\beta_{2}$ $\langle b_{4}^{\dagger 2}\rangle$, $\langle b_{5}^{2}\rangle$ $\langle\bm{\hat{B}_{2}^{2}}\rangle$ 1 1x4 1x4 1x3 1x3
$2\zeta$ $\langle b_{4}^{\dagger}b_{5}\rangle$ 2
$2\zeta-2\beta_{1}$ $\langle b_{2}^{2}\rangle$ $\langle\bm{\hat{B}_{3}^{2}}\rangle$ 1 1x1 1x1 1x1 1x1
$-2\zeta+2\beta_{1}$ $\langle b_{2}^{\dagger 2}\rangle$ $\langle\bm{\hat{B}_{4}^{2}}\rangle$ 1 1x1 1x1 1x1 1x1
$-2\zeta-2\beta_{1}$ $\langle b_{3}^{2}\rangle$ $\langle\bm{\hat{B}_{5}^{2}}\rangle$ 1 1x1 1x1 1x1 1x1
$2\zeta+2\beta_{1}$ $\langle b_{3}^{\dagger 2}\rangle$ $\langle\bm{\hat{B}_{6}^{2}}\rangle$ 1 1x1 1x1 1x1 1x1
0 $\langle b_{1}^{2}\rangle$, $\langle b_{1}^{\dagger 2}\rangle$ $\langle\bm{\hat{B}_{7}^{2}}\rangle$, $\langle\bm{\hat{B}_{8}^{2}}\rangle$ 1 4x1 4x1 3x1 3x1
$\langle b_{1}^{\dagger}b_{1}\rangle$ $\langle\bm{\hat{B}_{8}\hat{B}_{7}}\rangle$ 2
±$\pm 2\beta_{2}$ $\langle b_{4}^{\dagger}b_{5}^{\dagger}\rangle$, $\langle b_{4}b_{5}\rangle$ $\langle\bm{\hat{B}_{2}\hat{B}_{1}}\rangle$ 2 2x4 2x4 1x4 1x4
$\beta_{2}-\beta_{2}$ $\langle b_{4}^{\dagger}b_{4}\rangle$, $\langle b_{5}b_{5}^{\dagger}\rangle$ 2
$-\beta_{1}\mp\beta_{2}$ $\langle b_{4}b_{2}\rangle$, $\langle b_{5}^{\dagger}b_{2}\rangle$ $\langle\bm{\hat{B}_{1}\hat{B}_{3}}\rangle$ 2 4x1 4x1 2x1 2x1
$2\zeta-\beta_{1}\pm\beta_{2}$ $\langle b_{4}^{\dagger}b_{2}\rangle$, $\langle b_{5}b_{2}\rangle$ $\langle\bm{\hat{B}_{2}\hat{B}_{3}}\rangle$ 2 4x1 4x1 2x1 2x1
$-2\zeta+\beta_{1}\mp\beta_{2}$ $\langle b_{2}^{\dagger}b_{4}\rangle$, $\langle b_{2}^{\dagger}b_{5}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{4}\hat{B}_{1}}\rangle$ 2 4x1 4x1 2x1 2x1
$\beta_{1}\pm\beta_{2}$ $\langle b_{2}^{\dagger}b_{4}^{\dagger}\rangle$, $\langle b_{2}^{\dagger}b_{5}\rangle$ $\langle\bm{\hat{B}_{4}\hat{B}_{2}}\rangle$ 2 4x1 4x1 2x1 2x1
$-2\zeta-\beta_{1}\mp\beta_{2}$ $\langle b_{4}b_{3}\rangle$, $\langle b_{5}^{\dagger}b_{3}\rangle$ $\langle\bm{\hat{B}_{1}\hat{B}_{5}}\rangle$ 2 4x1 4x1 2x1 2x1
$-\beta_{1}\pm\beta_{2}$ $\langle b_{4}^{\dagger}b_{3}\rangle$, $\langle b_{5}b_{3}\rangle$ $\langle\bm{\hat{B}_{2}\hat{B}_{5}}\rangle$ 2 4x1 4x1 2x1 2x1
$\beta_{1}\mp\beta_{2}$ $\langle b_{3}^{\dagger}b_{4}\rangle$, $\langle b_{3}^{\dagger}b_{5}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{6}\hat{B}_{1}}\rangle$ 2 4x1 4x1 2x1 2x1
$2\zeta+\beta_{1}\pm\beta_{2}$ $\langle b_{3}^{\dagger}b_{4}^{\dagger}\rangle$, $\langle b_{3}^{\dagger}b_{5}\rangle$ $\langle\bm{\hat{B}_{6}\hat{B}_{2}}\rangle$ 2 4x1 4x1 2x1 2x1
$-\zeta\mp\beta_{2}$ $\langle b_{4}b_{1}\rangle$, $\langle b_{5}^{\dagger}b_{1}\rangle$ $\langle\bm{\hat{B}_{7}}\bm{\hat{B}_{1}}\rangle$ 2 4x2 4x2 2x2 2x2
$\langle b_{4}b_{1}^{\dagger}\rangle$, $\langle b_{5}^{\dagger}b_{1}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{8}}\bm{\hat{B}_{1}}\rangle$ 2
$\zeta\pm\beta_{2}$ $\langle b_{4}^{\dagger}b_{1}\rangle$, $\langle b_{5}b_{1}\rangle$ $\langle\bm{\hat{B}_{7}}\bm{\hat{B}_{2}}\rangle$ 2 4x2 4x2 2x2 2x2
$\langle b_{4}^{\dagger}b_{1}^{\dagger}\rangle$, $\langle b_{5}b_{1}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{8}}\bm{\hat{B}_{2}}\rangle$ 2
$\beta_{1}-\beta_{1}$ $\langle b_{2}^{\dagger}b_{2}\rangle$ $\langle\bm{\hat{B}_{4}\hat{B}_{3}}\rangle$ 2 2x1 2x1 1x1 1x1
$-2\beta_{1}$ $\langle b_{3}b_{2}\rangle$ $\langle\bm{\hat{B}_{5}\hat{B}_{3}}\rangle$ 2 2x1 2x1 1x1 1x1
$2\zeta$ $\langle b_{3}^{\dagger}b_{2}\rangle$ $\langle\bm{\hat{B}_{6}\hat{B}_{3}}\rangle$ 2 2x1 2x1 1x1 1x1
$\zeta-\beta_{1}$ $\langle b_{1}b_{2}\rangle$ $\langle\bm{\hat{B}_{7}}\bm{\hat{B}_{3}}\rangle$ 2 4x1 4x1 2x1 2x1
$\langle b_{1}^{\dagger}b_{2}\rangle$ $\langle\bm{\hat{B}_{8}}\bm{\hat{B}_{3}}\rangle$ 2
$-2\zeta$ $\langle b_{2}^{\dagger}b_{3}\rangle$ $\langle\bm{\hat{B}_{4}\hat{B}_{5}}\rangle$ 2 2x1 2x1 1x1 1x1
$2\beta_{1}$ $\langle b_{2}^{\dagger}b_{3}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{4}\hat{B}_{6}}\rangle$ 2 2x1 2x1 1x1 1x1
$-\zeta-\beta_{1}$ $\langle b_{2}^{\dagger}b_{1}\rangle$ $\langle\bm{\hat{B}_{4}}\bm{\hat{B}_{7}}\rangle$ 2 4x1 4x1 2x1 2x1
$\langle b_{2}^{\dagger}b_{1}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{4}}\bm{\hat{B}_{8}}\rangle$ 2
$\beta_{1}-\beta_{1}$ $\langle b_{3}^{\dagger}b_{3}\rangle$ $\langle\bm{\hat{B}_{6}\hat{B}_{5}}\rangle$ 2 2x1 2x1 1x1 1x1
$-\zeta-\beta_{1}$ $\langle b_{1}b_{3}\rangle$ $\langle\bm{\hat{B}_{7}}\bm{\hat{B}_{5}}\rangle$ 2 4x1 4x1 2x1 2x1
$\langle b_{1}^{\dagger}b_{3}\rangle$ $\langle\bm{\hat{B}_{8}}\bm{\hat{B}_{5}}\rangle$ 2
$\zeta+\beta_{1}$ $\langle b_{3}^{\dagger}b_{1}\rangle$ $\langle\bm{\hat{B}_{6}}\bm{\hat{B}_{7}}\rangle$ 2 4x1 4x1 2x1 2x1
$\langle b_{3}^{\dagger}b_{1}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{6}}\bm{\hat{B}_{8}}\rangle$ 2
@@ 0 $\langle b_{1}\rangle$ $\langle\bm{\hat{B}_{7}}\rangle$ 1 2x1 2x1 2x1 2x1
Table 6: Real and imaginary parts of the complex eigenfrequencies $\bm{\Lambda_{j}^{\rm r}}-i\bm{\Lambda_{j}^{\rm i}}$ of the matrix $\bm{M^{(5)}_{\rm p}}$, given in Eq. (81), for the five-mode pyramid bosonic system derived from the equations for the FOMs up to second order assuming $\beta_{2}=0$. For details, see the caption to Tab. 1. If $\beta_{1}=0$ instead of $\beta_{2}=0$, simple relabelling of $b_{j}$, $b_{j}^{\dagger}$ operators provides the corresponding table.
VI Conclusions
We have analyzed the dynamics of simple bosonic systems described by quadratic non-Hermitian Hamiltonians from the point of view of the occurrence of quantum exceptional, diabolical, and hybrid points. Non-Hermiticity of the considered systems, composed of from two to five coupled modes, originated in their damping and/or amplification, that are accompanied by the corresponding Langevin fluctuating forces to assure the physically consistent behavior. We have identified specific configurations defined by two-mode couplings and conditions for the damping and amplification rates of the modes at which the inherited quantum exceptional and diabolical points occur. Surprisingly, in these physically consistent models, we have found only second- and third-order inherited quantum exceptional points including their doubling due to second-order diabolical degeneracies. We have shown in general that the analyzed bosonic systems naturally exhibit these second-order diabolical degeneracies. The exceptional and diabolical degeneracies of inherited quantum hybrid points have then been used to construct higher-order degeneracies observed in the dynamics of second-order field-operator moments. The corresponding quantum exceptional and hybrid points are summarized in tables that demonstrate a rich structure of the evolution of the general-order field-operator moments.
The investigations have revealed the need for further looking for the bosonic systems with higher-order inherited quantum exceptional and hybrid points by weakening the requirements to the considered systems. In Ref. [50] we extend our investigations to bosonic systems with partial $\mathcal{PT}$-symmetry like dynamics (nonconventional $\mathcal{PT}$-symmetry) as well as non-Hermitian bosonic systems with unidirectional coupling.
VII Acknowledgements
The authors thank Ievgen I. Arkhipov for useful discussions. J.P. and K.T. acknowledge support by the project OP JAC CZ.02.01.01/00/22_008/0004596 of the Ministry of Education, Youth, and Sports of the Czech Republic. A.K.-K., G.Ch., and A.M. were supported by the Polish National Science Centre (NCN) under the Maestro Grant No. DEC-2019/34/A/ST2/00081.
A.1 General $n$-mode bosonic systems
$\Lambda_{2j-1}^{M^{(n)}}$ $=$ ${\lambda_{j}^{M^{(n)}}}({\xi}=\lambda^{\xi}_{1}),$
[…]
A.2 Three-mode bosonic system
A.3 Four-mode bosonic systems
In the four-mode system ($n=4$) in the linear configuration with different damping and/or amplification rates in neighbor modes and assuming $\alpha_{+}=0$ [$\alpha_{-}=0$], $\alpha_{\pm}=\sqrt{(3\pm\sqrt{5})\zeta^{2}/2-\gamma_{-}^{2})}=0$, we find a single QHP with second-order ED and DD. The eigenvalues $\Lambda_{1,2,3,4}^{M^{(4)}_{\rm l2}}$ [$\Lambda_{5,6,7,8}^{M^{(4)}_{\rm l2}}$] in Eq. (87) are equal. The eigenvectors in Eq. (97) fulfil: $\bm{Y_{1}^{M^{(4)}_{\rm l2}}}=\bm{Y_{3}^{M^{(4)}_{\rm l3}}}$ and $\bm{Y_{2}^{M^{(4)}_{\rm l2}}}=\bm{Y_{4}^{M^{(4)}_{\rm l2}}}$ [$\bm{Y_{5}^{M^{(4)}_{\rm l2}}}=\bm{Y_{7}^{M^{(4)}_{\rm l2}}}$ and $\bm{Y_{6}^{M^{(4)}_{\rm l2}}}=\bm{Y_{8}^{M^{(4)}_{\rm l2}}}$].
A.4 Five-mode bosonic systems

### v3

@@ 0 $\langle\hat{b}_{1}\rangle$ $\langle\bm{\hat{B}_{5}}\rangle$ 1 2x1 2x1 2x1 2x1
$\gamma_{+}$ $\zeta\mp\beta_{2}$ $\langle\hat{b}_{4}\rangle$, $\langle\hat{b}_{5}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{1}}\rangle$ 1 1x2 1x2 1x2 1x2
$\zeta\pm\beta_{2}$ $\langle\hat{b}_{4}^{\dagger}\rangle$, $\langle\hat{b}_{5}\rangle$ $\langle\bm{\hat{B}_{2}}\rangle$ 1 1x2 1x2 1x2 1x2
$\zeta-\beta_{1}$ $\langle\hat{b}_{2}\rangle$ $\langle\bm{\hat{B}_{3}}\rangle$ 1 1x1 1x1 1x1 1x1
$\zeta+\beta_{1}$ $\langle\hat{b}_{2}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{4}}\rangle$ 1 1x1 1x1 1x1 1x1
$\zeta-\beta_{1}$ $\langle\hat{b}_{3}\rangle$ $\langle\bm{\hat{B}_{5}}\rangle$ 1 1x1 1x1 1x1 1x1
$\zeta+\beta_{1}$ $\langle\hat{b}_{3}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{6}}\rangle$ 1 1x1 1x1 1x1 1x1
0 $\langle\hat{b}_{1}\rangle$ $\langle\bm{\hat{B}_{7}}\rangle$ 1 2x1 2x1 2x1 2x1
$\langle\hat{b}_{1}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{8}}\rangle$ 1
$2\gamma_{+}$ $2\zeta\mp 2\beta_{2}$ $\langle\hat{b}_{4}^{2}\rangle$, $\langle\hat{b}_{5}^{\dagger 2}\rangle$ $\langle\bm{\hat{B}_{1}^{2}}\rangle$ 1 1x4 1x4 1x3 1x3
$2\zeta$ $\langle\hat{b}_{5}^{\dagger}\hat{b}_{4}\rangle$ 2
$2\zeta\pm 2\beta_{2}$ $\langle\hat{b}_{4}^{\dagger 2}\rangle$, $\langle\hat{b}_{5}^{2}\rangle$ $\langle\bm{\hat{B}_{2}^{2}}\rangle$ 1 1x4 1x4 1x3 1x3
$2\zeta$ $\langle\hat{b}_{4}^{\dagger}\hat{b}_{5}\rangle$ 2
$2\zeta-2\beta_{1}$ $\langle\hat{b}_{2}^{2}\rangle$ $\langle\bm{\hat{B}_{3}^{2}}\rangle$ 1 1x1 1x1 1x1 1x1
$2\zeta+2\beta_{1}$ $\langle\hat{b}_{2}^{\dagger 2}\rangle$ $\langle\bm{\hat{B}_{4}^{2}}\rangle$ 1 1x1 1x1 1x1 1x1
$2\zeta-2\beta_{1}$ $\langle\hat{b}_{3}^{2}\rangle$ $\langle\bm{\hat{B}_{5}^{2}}\rangle$ 1 1x1 1x1 1x1 1x1
$2\zeta+2\beta_{1}$ $\langle\hat{b}_{3}^{\dagger 2}\rangle$ $\langle\bm{\hat{B}_{6}^{2}}\rangle$ 1 1x1 1x1 1x1 1x1
0 $\langle\hat{b}_{1}^{2}\rangle$, $\langle\hat{b}_{1}^{\dagger 2}\rangle$ $\langle\bm{\hat{B}_{7}^{2}}\rangle$, $\langle\bm{\hat{B}_{8}^{2}}\rangle$ 1 4x1 4x1 3x1 3x1
$\langle\hat{b}_{1}^{\dagger}\hat{b}_{1}\rangle$ $\langle\bm{\hat{B}_{8}\hat{B}_{7}}\rangle$ 2
$\pm 2\beta_{2}$ $\langle\hat{b}_{4}^{\dagger}\hat{b}_{5}^{\dagger}\rangle$, $\langle\hat{b}_{4}\hat{b}_{5}\rangle$ $\langle\bm{\hat{B}_{2}\hat{B}_{1}}\rangle$ 2 2x4 2x4 1x4 1x4
$\beta_{2}-\beta_{2}$ $\langle\hat{b}_{4}^{\dagger}\hat{b}_{4}\rangle$, $\langle\hat{b}_{5}\hat{b}_{5}^{\dagger}\rangle$ 2
$\beta_{1}\mp\beta_{2}$ $\langle\hat{b}_{4}\hat{b}_{2}\rangle$, $\langle\hat{b}_{5}^{\dagger}\hat{b}_{2}\rangle$ $\langle\bm{\hat{B}_{1}\hat{B}_{3}}\rangle$ 2 4x1 4x1 2x1 2x1
$2\zeta-\beta_{1}\pm\beta_{2}$ $\langle\hat{b}_{4}^{\dagger}\hat{b}_{2}\rangle$, $\langle\hat{b}_{5}\hat{b}_{2}\rangle$ $\langle\bm{\hat{B}_{2}\hat{B}_{3}}\rangle$ 2 4x1 4x1 2x1 2x1
$2\zeta+\beta_{1}\mp\beta_{2}$ $\langle\hat{b}_{2}^{\dagger}\hat{b}_{4}\rangle$, $\langle\hat{b}_{2}^{\dagger}\hat{b}_{5}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{4}\hat{B}_{1}}\rangle$ 2 4x1 4x1 2x1 2x1
$\beta_{1}\pm\beta_{2}$ $\langle\hat{b}_{2}^{\dagger}\hat{b}_{4}^{\dagger}\rangle$, $\langle\hat{b}_{2}^{\dagger}\hat{b}_{5}\rangle$ $\langle\bm{\hat{B}_{4}\hat{B}_{2}}\rangle$ 2 4x1 4x1 2x1 2x1
$2\zeta-\beta_{1}\mp\beta_{2}$ $\langle\hat{b}_{4}\hat{b}_{3}\rangle$, $\langle\hat{b}_{5}^{\dagger}\hat{b}_{3}\rangle$ $\langle\bm{\hat{B}_{1}\hat{B}_{5}}\rangle$ 2 4x1 4x1 2x1 2x1
$\beta_{1}\pm\beta_{2}$ $\langle\hat{b}_{4}^{\dagger}\hat{b}_{3}\rangle$, $\langle\hat{b}_{5}\hat{b}_{3}\rangle$ $\langle\bm{\hat{B}_{2}\hat{B}_{5}}\rangle$ 2 4x1 4x1 2x1 2x1
$\beta_{1}\mp\beta_{2}$ $\langle\hat{b}_{3}^{\dagger}\hat{b}_{4}\rangle$, $\langle\hat{b}_{3}^{\dagger}\hat{b}_{5}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{6}\hat{B}_{1}}\rangle$ 2 4x1 4x1 2x1 2x1
$2\zeta+\beta_{1}\pm\beta_{2}$ $\langle\hat{b}_{3}^{\dagger}\hat{b}_{4}^{\dagger}\rangle$, $\langle\hat{b}_{3}^{\dagger}\hat{b}_{5}\rangle$ $\langle\bm{\hat{B}_{6}\hat{B}_{2}}\rangle$ 2 4x1 4x1 2x1 2x1
$\zeta\mp\beta_{2}$ $\langle\hat{b}_{4}\hat{b}_{1}\rangle$, $\langle\hat{b}_{5}^{\dagger}\hat{b}_{1}\rangle$ $\langle\bm{\hat{B}_{7}}\bm{\hat{B}_{1}}\rangle$ 2 4x2 4x2 2x2 2x2
$\langle\hat{b}_{4}\hat{b}_{1}^{\dagger}\rangle$, $\langle\hat{b}_{5}^{\dagger}\hat{b}_{1}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{8}}\bm{\hat{B}_{1}}\rangle$ 2
$\zeta\pm\beta_{2}$ $\langle\hat{b}_{4}^{\dagger}\hat{b}_{1}\rangle$, $\langle\hat{b}_{5}\hat{b}_{1}\rangle$ $\langle\bm{\hat{B}_{7}}\bm{\hat{B}_{2}}\rangle$ 2 4x2 4x2 2x2 2x2
$\langle\hat{b}_{4}^{\dagger}\hat{b}_{1}^{\dagger}\rangle$, $\langle\hat{b}_{5}\hat{b}_{1}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{8}}\bm{\hat{B}_{2}}\rangle$ 2
$\beta_{1}-\beta_{1}$ $\langle\hat{b}_{2}^{\dagger}\hat{b}_{2}\rangle$ $\langle\bm{\hat{B}_{4}\hat{B}_{3}}\rangle$ 2 2x1 2x1 1x1 1x1
$2\beta_{1}$ $\langle\hat{b}_{3}\hat{b}_{2}\rangle$ $\langle\bm{\hat{B}_{5}\hat{B}_{3}}\rangle$ 2 2x1 2x1 1x1 1x1
$2\zeta$ $\langle\hat{b}_{3}^{\dagger}\hat{b}_{2}\rangle$ $\langle\bm{\hat{B}_{6}\hat{B}_{3}}\rangle$ 2 2x1 2x1 1x1 1x1
$\zeta-\beta_{1}$ $\langle\hat{b}_{1}\hat{b}_{2}\rangle$ $\langle\bm{\hat{B}_{7}}\bm{\hat{B}_{3}}\rangle$ 2 4x1 4x1 2x1 2x1
$\langle\hat{b}_{1}^{\dagger}\hat{b}_{2}\rangle$ $\langle\bm{\hat{B}_{8}}\bm{\hat{B}_{3}}\rangle$ 2
$2\zeta$ $\langle\hat{b}_{2}^{\dagger}\hat{b}_{3}\rangle$ $\langle\bm{\hat{B}_{4}\hat{B}_{5}}\rangle$ 2 2x1 2x1 1x1 1x1
$2\beta_{1}$ $\langle\hat{b}_{2}^{\dagger}\hat{b}_{3}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{4}\hat{B}_{6}}\rangle$ 2 2x1 2x1 1x1 1x1
$\zeta-\beta_{1}$ $\langle\hat{b}_{2}^{\dagger}\hat{b}_{1}\rangle$ $\langle\bm{\hat{B}_{4}}\bm{\hat{B}_{7}}\rangle$ 2 4x1 4x1 2x1 2x1
$\langle\hat{b}_{2}^{\dagger}\hat{b}_{1}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{4}}\bm{\hat{B}_{8}}\rangle$ 2
$\beta_{1}-\beta_{1}$ $\langle\hat{b}_{3}^{\dagger}\hat{b}_{3}\rangle$ $\langle\bm{\hat{B}_{6}\hat{B}_{5}}\rangle$ 2 2x1 2x1 1x1 1x1
$\zeta-\beta_{1}$ $\langle\hat{b}_{1}\hat{b}_{3}\rangle$ $\langle\bm{\hat{B}_{7}}\bm{\hat{B}_{5}}\rangle$ 2 4x1 4x1 2x1 2x1
$\langle\hat{b}_{1}^{\dagger}\hat{b}_{3}\rangle$ $\langle\bm{\hat{B}_{8}}\bm{\hat{B}_{5}}\rangle$ 2
$\zeta+\beta_{1}$ $\langle\hat{b}_{3}^{\dagger}\hat{b}_{1}\rangle$ $\langle\bm{\hat{B}_{6}}\bm{\hat{B}_{7}}\rangle$ 2 4x1 4x1 2x1 2x1
$\langle\hat{b}_{3}^{\dagger}\hat{b}_{1}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{6}}\bm{\hat{B}_{8}}\rangle$ 2
@@ 0 $\langle\hat{b}_{1}\rangle$ $\langle\bm{\hat{B}_{7}}\rangle$ 1 2x1 2x1 2x1 2x1
Table 6: Real and imaginary parts of the complex eigenfrequencies $\bm{\Lambda_{j}^{\rm r}}-i\bm{\Lambda_{j}^{\rm i}}$ of the matrix $\bm{M^{(5)}_{\rm p}}$, given in Eq. (86), for the five-mode pyramid bosonic system derived from the equations for the FOMs up to second order assuming $\beta_{2}=0$. If $\beta_{1}=0$ instead of $\beta_{2}=0$, simple relabelling of $\hat{b}_{j}$, $\hat{b}_{j}^{\dagger}$ operators provides the corresponding table. We have $\bm{\hat{B}_{1}}\equiv[\hat{b}_{4},\hat{b}_{5}^{\dagger}]$, $\bm{\hat{B}_{2}}=\bm{\hat{B}_{1}}^{\dagger}$, $\bm{\hat{B}_{3}}\equiv[\hat{b}_{2}]$, $\bm{\hat{B}_{4}}=\bm{\hat{B}_{3}}^{\dagger}$, $\bm{\hat{B}_{5}}\equiv[\hat{b}_{3}]$, $\bm{\hat{B}_{6}}=\bm{\hat{B}_{5}}^{\dagger}$, $\bm{\hat{B}_{7}}\equiv[\hat{b}_{1}]$, $\bm{\hat{B}_{8}}=\bm{\hat{B}_{7}}^{\dagger}$, and more details are given in the caption to Tab. 1.
Finally, for the five-mode bosonic system in the pyramid configuration, described by the dynamical matrix ${\bm{M}}_{\rm p}^{(5)}$ given in Eq. (86) with damping and/or amplification rates satisfying the conditions in Eqs. (87) and (88), the QEPs and QHPs identified in the dynamics of first- and second-order FOMs are systematically summarized in Tab. 6. Genuine QEPs with second-, third-, and fourth-order EDs appear in the spectral structure of second-order FOMs, though with relatively low DDs. This reflects the limited presence of DDs in the first-order FOM spectra, where only one eigenfrequency exhibits a second-order DD.